OpenStudy (anonymous):

Which of the following aqueous solutions will have the lowest vapor pressure at 25°C?

OpenStudy (anonymous):

1.5 M C6H12O6 1.5 M LiNO3 1.0 M Al2O3 1.0 M CaF2

OpenStudy (anonymous):

1. How many grams of a stock solution that is 87.5 percent H2SO4 by mass would be needed to make 275 grams of a 55.0 percent by mass solution? Show all of the work needed to solve this problem. Your final solution contains 55% by mass of H2SO4. You have 275 g of that solution, so it contains 0.55 x 275 g of H2SO4 i.e. 151.25 g H2SO4. The stock solution contains 87.5% H2SO4. So 100 g of that solution contains 87.5 g of H2SO4. You need 151.25 g of H2SO4, so now it’s just a simple proportion question: if you call the mass of stock solution m, then … m/100 = 151.25 / 87.5 so m = 100 (151.25 / 87.5) = 173.8 g ______________________________________… 2. Describe, in detail, how to make 3.0 liters of a 0.98 molar HCl solution from a 10.5 molar HCl stock solution. (Your answer is right. But the correct unit for molar concentration is mol/L. The old symbol M is not part of the S.I. unit system, and was abandoned in about 1972. ) ______________________________________… 3. How many grams of zinc metal will react completely with 8.2 liters of 3.5 M HCl? Show all of the work needed to solve this problem. Zn (s) + 2 HCl (aq) ---------> ZnCl2 (aq) + H2 (g) I don't understand it. : ( From the equation, you can see that 1 mol Zn will react with 2 mol HCl. So first you have to calculate how many moles of HCl are available, then work out how many moles of zinc will react with that acid, then convert the moles of zinc to grams. Using the usual symbols [ n = no. of moles (mol), c = concentration (mol/L), V = volume (L) ] … n = cV = (3.5 mol/L)(8.2 L) = 28.7 mol From the equation, 2 mol HCl reacts with 1 mol Zn. So 28.7 mol HCl will react with 14.35 mol Zn. Now convert 14.35 mol Zn to grams: mass (g) = molar mass (g/mol) x moles (mol) m = (65.4 g/mol)(14.35 mol) = 938.5 g ______________________________________… 4. How many grams of iron metal do you expect to be produced when 325 grams of an 87.5 percent by mass iron (II) nitrate solution reacts with excess aluminum metal? Show all of the work needed to solve this problem. 2 Al (s) + 3 Fe(NO3)2 (aq) ----------> 3 Fe (s) + 2 Al(NO3)2 (aq) Same with this one. This is similar to the previous question, with the added annoyance of having to deal with this impossible 87.5% solution. (You simply can’t create such a solution: iron II nitrate isn’t soluble enough. The maximum possible concentration is about 45%.) Anyway … Fom the equation you can see that 3 mol of iron II nitrate will produce 3 mol of iron. You need to figure out how many moles of iron II nitrate you have available: that will instantly tell you how many moles of iron will be produced: and then you can convert that amount to grams. The solution is 87.5%. You’re using 325 g of it, so the mass of iron II nitrate available is 87.5% of 325 g, which comes to 284.4 g. The molar mass of iron II nitrate is 179.9 g/mol: the number of moles of iron II nitrate is calculated using moles of iron II nitrate = mass (g) / molar mass (g/mol) = 284.4 g / 179.9 g/mol = 1.58 mol Since it’s a 1:1 reaction, you’ll end up with 1.58 moles of Fe. Convert that to grams: mass of Fe = moles x molar mass = (1.58 mol)(55.9 g/mol) = 88.3 g ______________________________________… 5. is correct 6. is correct. 7. is not correct. You need to use …. number of moles (mol) = molar concentration (mol/L) x volume (L) ______________________________________… The other questions I’ll leave to you. The principle is simple. When you dissolve a material (a ‘non-volatile solute’) in water the dissolved particles (molecules or ions) cling to the water molecules, and they also keep the water molecules apart from each other. The clinging makes it harder to evaporate or boil the solution. The forcing-apart makes it harder for the water molecules to get close enough to stick together and freeze out as the temperature drops. The more ‘particles’ of solute are in the solution, the more exaggerated these effects become. When you dissolve an ionic material in water it dissociates, so a mole of NaCl will give 2 moles of ions, while a mole of CaCl2 will give 3 moles of ions, and so on. (Methane, CH4, is NOT a non-volatile solute, so it doesn’t have the same behaviour as the ionic materials in question 9.) I typed this rather quickly, so you should check my calculations to make sure they’re correct. I offer no guarantee of correctness! Hope this is of some use to you. :)

OpenStudy (anonymous):

had this ? in a report from a friend hope it helps

OpenStudy (anonymous):

copied it from her desk top

OpenStudy (anonymous):

did it help u

OpenStudy (anonymous):

No. :-/ It was rather confusing! D:

OpenStudy (anonymous):

sorry D2 * 3 = 6 larger than others C look up colligative properties 3. 0C – 25.5 g/MW of glocuse * 1000/398 *1.86C 4. 100C + 2.5/198 *1000 *0.51C

OpenStudy (anonymous):

Vapor pressure is inversely related to intermolecular forces. So if you want to find which one will have the greatest vapor pressure, then draw out their molecular shapes (3-D) and then you take the ones with the least intermolecular forces.

OpenStudy (anonymous):

i am of no help sorry

OpenStudy (anonymous):

haha yeah you copied that from a site. :P

OpenStudy (anonymous):

nine you gott me!!sorry mad you sounded desperate and i had to do some thing :(